OpenStudy (anonymous):

Find an angle between 0 and 2π that is coterminal with the given angle. Give answer in terms of π. -2pi/3 . ? in radians

OpenStudy (anonymous):

all you have to do is add or subtract 2pi or some other amount of 2pi for each one in this one, add 2pi to -2pi/3

OpenStudy (anonymous):

How do I know if I have add or subtract ? ==

OpenStudy (anonymous):

is - 2pi/3 in the interval from 0 to 2pi? if it's not then, add or subtract 2pi so the angle is between 0 and 2pi

OpenStudy (anonymous):

\[-\frac{2\pi}{3} + 2\pi = -\frac{2\pi}{3} + \frac{6\pi}{3}\]

OpenStudy (anonymous):

So confuse ! So how do I know when I have to ADD OR SUBTRACT ?